# Geometric Model Selection for Latent Space Network Models: Hypothesis Testing via Multidimensional Scaling and Resampling Techniques

### v1

@@ 1 Introduction
Current approaches to determining which geometry underlies a particular observed network include: (i) comparing the fit of multi-dimensional scaling (MDS) of geodesic network distances across different geometries, as in Papamichalis et al. [2022]; (ii) borrowing tools from metric spaces, such as $\delta$-hyperbolicity (Kennedy et al. [2013];Narayan and Saniee [2011]) or Van Der Hoorn et al. [2021]; and (iii) curvature estimation, via a clique-based latent distance estimator (Lubold et al. [2022]; Wilkins-Reeves and McCormick [2024]).
@@
In Section 2, we review MDS in hyperbolic and Euclidean spaces and show how comparisons of standard goodness-of-fit measures, stress and strain, can guide geometry choice. In Section 3, we introduce methods for quantifying uncertainty for MDS goodness-of-fit metrics, reviewing the permutation test and extending the bootstrap $J$-test to the network setting. In Section 4, we describe the latent space models in both hyperbolic and Euclidean space and demonstrate the efficiency and accuracy of the methods through simulations. Section 5 demonstrates the methods on several real-world datasets.
@@ 2 Multi-dimensional scaling in the network setting
Moreover, only connected networks are considered in this study; that is, all the elements in the geodesic distance matrix $\mathbf{D}(\mathbf{Y})$ are non-negative finite numbers.
Classical MDS ([Borg and Groenen, 2007, chap 12]), also known as principal coordinate analysis, uses observed dissimilarities to embed data in a lower-dimensional metric space. Given a dissimilarity matrix $\boldsymbol{\Delta}$, the goal is to find a coordinate matrix $\mathbf{X}_{\mathbb{M}}$ in a low-dimensional manifold $\mathbb{M}$ that minimizes a target function which measures stress or strain. Stress serves as a goodness-of-fit measure for MDS. Although multiple definitions of stress have been proposed, we use the standard definition given by,
$S_{\mathbb{M}}(\mathbf{Y})=\sqrt{\sum_{i,j}(\delta_{ij}-\hat{\delta}_{ij})^{2}}$
where $\mathbf{Y}$ is the adjacency matrix of the observed network, $\delta_{ij}$ is the measure of dissimilarity (e.g. geodesic distance, or shortest path lengths) between nodes $i$ and $j$, and $\hat{\delta}_{ij}$ is the distance between points $n_{i}$ and $n_{j}$ calculated in manifold $\mathbb{M}$ based on the coordinate matrix $\mathbf{X}_{\mathbb{M}}$ recovered by MDS.
2.1 Method 1: Observed Stress Difference
In Papamichalis et al. [2022], they compare the quality of multidimensional scaling (MDS) across different underlying geometries. The stress from classical and hyperbolic MDS are compared directly. The geodesic distance matrix is used as the dissimilarity matrix and if $S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y})<0$, then the hyperbolic space is preferred as the underlying latent space.
2.2 Method 2: Permutation Test
In the literature, both permutation tests and bootstrapping have been used to measure uncertainty associated with MDS goodness of fit metrics Mair et al. [2016], Mair et al. [2022], Farine and Carter [2022]
@@ 2.2 Method 2: Permutation Test
$\mathbf{H_{0}}:$ Stress/configuration are obtained from a random permutation of dissimilarities.
$\mathbf{H_{A}}:$ Stress/configuration are obtained from something other than a random permutation of dissimilarities.
This null hypothesis is weakly informative and states that dissimilarities are exchangeable and that there is no structure beyond random chance. Translating these hypotheses to the comparison between MDS approaches and the latent space network setting, the null and alternative hypothesis are given as
@@
S1: In the case of directly observed dissimilarities, the elements of $\boldsymbol{\Delta}$ can be permuted.
S2: For derived dissimilarities, they proposed a strategy for systematic column-wise permutations of the raw data.
@@
This method is described in the following:
Algorithm 1 Method 2: Permutation Test
@@
4:   Apply classical MDS and hydra on $\mathbf{D}(\tilde{\mathbf{Y}}_{i})$
@@
2.3 Method 3: Bootstrapping
Bootstrapping Networks
Levin and Levina [2021]
@@ 2.3 Method 3: Bootstrapping
provide two methods to bootstrap a network. The first method is designed to generate bootstrap replicates of U-statistics, which does not apply to our case. For the second method, they generate bootstrap replicates of whole networks and then evaluate network quantities. They work with the random dot product graph(RDPG) in which the latent positions $X_{i}$ are i.i.d. with inner product distribution $F$. In order to obtain bootstrapping replicates, they apply adjacency spectral embedding(ASE) on the adjacency matrix $A$, and get estimated latent positions $\hat{X}_{i}$. Then, let $\hat{F}_{n}$ denote the empirical distribution of $\hat{X}_{i}$ and bootstrap sample $X_{i}^{*}$ are drawn i.i.d. from $\hat{F}_{n}$. Bootstrap samples generated in this way have adjacency matrices similar to the observed one.
@@
In our setting, we apply MDS to the geodesic distance matrix and get a coordinate matrix, but this matrix is not the latent positions. To generate bootstrap samples with a geodesic distance matrix that have a distribution similar to the observed network, we try to associate the geodesic distance with the distance between two nodes. In this new bootstrapping method, we estimate the distribution of the distance between two nodes conditioned on their observed geodesic distance and generate the distance matrix based on it. Then, generate a new bootstrap adjacency matrix.
Distribution of geodesic distance between two nodes given their positions
Fronczak et al. [2004]
@@
discuss an analytic solution for the average path length in a random graph, within which they provide a method to find the probability that the geodesic distance of $n_{i}$ and $n_{j}$ is exactly $k$. Rastelli et al. [2016] uses the same reasoning and finds explicit expression under Gaussian LPM. To summarize this method:
@@
First, define $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as the probability of at least one walk of length $k$ between $n_{i}$ and $n_{j}$. A walk is called a path only when all the nodes it passes through are distinct, while a walk does not require all its nodes to be distinct. In other words, $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ could also be thought of as the probability that the geodesic distance between these two nodes is not larger than $k$.
Then, define $\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as the probability that the geodesic distance between $n_{i}$ and $n_{j}$ is exactly $k$. Given the definition of $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$, we have:
@@
We further define $\xi_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ to be the probability that there exits a walk of length $k$ from $n_{i}$ to $n_{j}$ that passes through nodes $\{n_{i},n_{A_{m}^{1}},\dots,n_{A_{m}^{k-1}},n_{j}\}$, and we could refer to this walk of $k+1$ nodes event $A_{m}$. By the definition of a walk, we do not require $n_{A_{m}^{1}},\dots,n_{A_{m}^{k-1}}$ to be distinct. The number of such events is about $N^{k-1}$. Let $A=\cup_{m=1}^{N^{k-1}}A_{m}$, where $A$ is the event that at least one walk of length $k$ between $n_{i}$ and $n_{j}$. By Lemma 1 of Fronczak et al. [2004], we have:
@@
As suggested in Rastelli et al. [2016], when a network is given, setting $\gamma=1$, $\phi$ and $\tau$ could be found through ad-hoc method that matches the observed and theoretical value of average degree and clustering coefficient through equation (15) and (16).
Distribution of distance between nodes given their geodesic distance
By definition, we could write $\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as conditional probability, that is,
$\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|\mathbf{z}_{i},\mathbf{z}_{j})$ (8)
Based on the Gaussian LPM (14), we know that the probability of having an edge depend only on $d_{ij}=d_{\mathbb{R}^{d}}(\mathbf{z}_{i},\mathbf{z}_{j})$. Thus, we could also rewrite equation (8)
$\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|d_{ij})$ (9)
Then, the distribution of distance between nodes given their geodesic distance could be found through:
@@
When $d=2$ and $\gamma$ set to be $1$,
@@
Davidson-MacKinnon $J$-test
$J$
-test is a classical way to compare two nonnested regression models, and the core idea is that if model $M_{2}$ captures the missing structure of model $M_{1}$, then $M_{2}$’s fitted value should add explanatory power. In Section 5 of Davidson and MacKinnon [2002], they discuss the use of parametric bootstrap in $J$ test, with hypotheses:
$H_{0}$: The competing model does not provide additional explanatory power beyond the current model.
$H_{1}$: The competing model provides additional explanatory power.
To start with, they find $J$ statistic for the observed data set, denoted by $\hat{J}$. Then, to perform the bootstrapping, they fit the data into one of the two models and find the parameter estimates. These estimated parameters could later be used in the bootstrap data generating process. Once they have $B$ bootstrap samples, $J$ statistics are calculated again for these samples using the same method as the observed data set and are denoted by $J_{j}^{*},j=1,\dots,B$. Finally, the p-value for this test is given by
@@
In the case of the latent space model, we adopt this bootstrapping method of the Davidson-MacKinnon $J$ test. Similar to the Davidson-MacKinnon $J$-test, we are comparing two models: GLPM and hyperbolic geometry model. The hypotheses are:
$H_{0}$: The hyperbolic model does not provide additional explanatory power beyond the euclidean model.
$H_{1}$: The hyperbolic model provides additional explanatory power.
We first found the stress difference between classical MDS and hydra. Then, just as they estimate parameter to generate bootstrap data, by using the distribution of distance between nodes given geodesic distance, we could generate bootstrap networks according to GLPM. Then, calculate stress differences for these bootstrap networks and find the p-value. This method is described in the following:
Algorithm 2 Method 3: Bootstrapping
@@ 3 Simulations
We want to show the efficiency of the methods mentioned in the previous section. To achieve this goal, we simulate networks in Euclidean and hyperbolic latent spaces, using the models described in the following sections.
@@ 3.1 Latent space models
Modeling assumptions for latent space model are as follows:
@@ 3.1.1 Euclidean Geometry
Most commonly mentioned latent space models for networks that use Euclidean geometry as underlying latent space is the ”Distance Models” proposed in Hoff et al. [2002]. Later, Rastelli et al. [2016] proposed the Gaussian latent position model(Gaussian LPM, GLPM), in which a non-normalized Gaussian density is used for edge instead of logistic link function in Distance Models. We will use the Gaussian LPM as the Euclidean latent space models, and it’s given by:
Rastelli et al. [2016] has also shown that both expected average degree and clustering coefficient have explicit forms for Gaussian LPM, and are given by:

### v2

@@ 1 Introduction
Current approaches to determining which geometry underlies a particular observed network include: (i) comparing the fit of multi-dimensional scaling (MDS) of geodesic network distances across different geometries, as in Papamichalis et al. [2022]; (ii) borrowing tools from metric spaces, such as $\delta$-hyperbolicity (Kennedy et al. [2013]; Narayan and Saniee [2011]) or Ollivier-Ricci curvature (Van Der Hoorn et al. [2021]); and (iii) curvature estimation, via a clique-based latent distance estimator (Lubold et al. [2022]; Wilkins-Reeves and McCormick [2024]).
@@
In Section 2, we review MDS in hyperbolic and Euclidean spaces, show how comparisons of standard goodness-of-fit measures, stress and strain, can guide geometry choice, and discuss how permutation tests for quantifying MDS uncertainty can be adapted to the network setting. In Section 3, we introduce methods for quantifying uncertainty for MDS goodness-of-fit metrics, reviewing the permutation test and extending the bootstrap $J$-test to the network setting. In Section 4, we describe the latent space models in both hyperbolic and Euclidean space and demonstrate the efficiency and accuracy of the methods through simulations. Section 5 demonstrates the methods on several real-world datasets.
@@ 2 Multi-dimensional scaling in the network setting
Moreover, only connected networks are considered in this study; that is, all the elements $\delta_{ij}$ in the geodesic distance matrix $\mathbf{D}(\mathbf{Y})=(\delta_{ij})_{1\leq i,j\leq N}$ are non-negative finite numbers, where $\delta_{ij}$ is the shortest path length between $n_{i}$ and $n_{j}$.
Classical MDS ([Borg and Groenen, 2007, chap 12]), also known as principal coordinate analysis, uses observed dissimilarities to embed data in a lower-dimensional metric space. Given a dissimilarity matrix $\boldsymbol{\Delta}$, the goal is to find a coordinate matrix $\mathbf{X}_{\mathbb{M}}$ in a low-dimensional manifold $\mathbb{M}$ that minimizes a target function which measures stress or strain. Stress serves as a goodness-of-fit measure for MDS. Although multiple definitions of stress have been proposed, we use the definition given by Keller-Ressel and Nargang [2019],
$S_{\mathbb{M}}(\mathbf{Y})=\sqrt{\sum_{i,j}(\delta_{ij}-\hat{d}^{\mathbb{M}}_{ij})^{2}}$
where $\mathbf{Y}$ is the adjacency matrix of the observed network, $\delta_{ij}$ is the geodesic distance or shortest path lengths between nodes $i$ and $j$, which serves as a measure of dissimilarity in this study, and $\hat{d}^{\mathbb{M}}_{ij}$ is the distance between points $x_{i}$ and $x_{j}$ calculated in manifold $\mathbb{M}$ based on the coordinate matrix $\mathbf{X}_{\mathbb{M}}$ recovered by MDS.
In the following subsections, we briefly review how stress has been used to identify underlying latent network geometry and whether methods for quantifying uncertainty associated with MDS can be directly translated to the network setting.
2.1 Observed Stress Difference
In Papamichalis et al. [2022], they compare the quality of MDS across different underlying geometries. The stress from classical and hyperbolic MDS are compared directly. The geodesic distance matrix is used as the dissimilarity matrix and if $S_{\mathbb{H}^{2}}(\mathbf{Y})-S_{\mathbb{R}^{2}}(\mathbf{Y})<0$, then the hyperbolic space is preferred as the underlying latent space.
2.2 Permutation Tests
In the literature, both permutation tests and bootstrapping have been used to measure uncertainty associated with MDS goodness of fit metrics [Mair et al., 2016, 2022, Farine and Carter, 2022].
@@ 2.2 Permutation Tests
$\mathbf{H_{0}}$ :
Stress/configuration are obtained from a random permutation of dissimilarities.
$\mathbf{H_{A}}$ :
Stress/configuration are obtained from something other than a random permutation of dissimilarities.
This null hypothesis is weakly informative and states that dissimilarities are exchangeable and that there is no structure beyond random chance. Translating these hypotheses to the comparison between MDS approaches and the latent space network setting, the null and alternative hypotheses are given as
@@
S1 :
In the case of directly observed dissimilarities, the elements of $\boldsymbol{\Delta}$ can be permuted.
S2 :
For derived dissimilarities, they proposed a strategy for systematic column-wise permutations of the raw data.
@@
Algorithm 1 Permutation Test: Density-preserving permuted networks
@@
4:   Apply classical and hyperbolic MDS on $\mathbf{D}(\tilde{\mathbf{Y}}_{i})$
@@
3 Bootstrapped networks for MDS
The permuted $\tilde{\mathbf{Y}}$ form a nonparametric bootstrap distribution of networks. While preserving edge density, the nonparametric bootstrap distribution will fail to preserve other features of the underlying latent network geometry. For example, if two nodes have true latent positions which are close, they should be tied frequently in our bootstrap distribution. Rather than permuting the edges at random, we consider whether MDS provides information about the underlying latent distances between nodes.
@@ 3 Bootstrapped networks for MDS
Levin and Levina [2021] provide two methods to bootstrap networks that leverage estimated underlying latent structure. The first method is designed to generate bootstrap replicates of U-statistics, which does not apply to our case. For the second method, they generate bootstrap replicates of whole networks and then evaluate network quantities. They work with the random dot product graph (RDPG) in which the latent positions $X_{i}$ are i.i.d. with inner product distribution $F$. In order to obtain bootstrapping replicates, they apply adjacency spectral embedding (ASE) on the adjacency matrix $A$, and get estimated latent positions $\hat{X}_{i}$. Then, let $\hat{F}_{n}$ denote the empirical distribution of $\hat{X}_{i}$ and bootstrap samples, $X_{i}^{*}$, are drawn i.i.d. from $\hat{F}_{n}$. Bootstrap samples generated in this way have adjacency matrices similar to the observed one.
@@
In our setting, we apply MDS to the geodesic distance matrix, $\mathbf{D}(\mathbf{Y})$, and get a coordinate matrix, $X_{\mathbb{M}}$, recovered by MDS. It is tempting to use $X_{\mathbb{M}}$ in place of the ASE-estimated latent positions in [Levin and Levina, 2021]’s approach and draw a bootstrap replicate network based on the empirical distribution of $X_{\mathbb{M}}$. However, this matrix does not contain estimates of the true latent positions, $\boldsymbol{Z}$. Recall that the latent space network model and the RDPG assume that the likelihood of a tie between a pair of nodes is proportional to their latent distance. However, the pairwise distances induced by $X_{\mathbb{M}}$ correspond to geodesic distances on $Y$, and not tie presence directly. To generate bootstrap network samples with a geodesic distance matrix that have a distribution similar to the observed network, we will associate the geodesic distance with the underlying latent distance between two nodes. In this new bootstrapping method, we estimate the distribution of the distance between two nodes conditioned on their observed geodesic distance and generate the corresponding distance matrix. Then, generate a new bootstrap adjacency matrix.
3.1 Geodesic distances given latent positions
@@
Fronczak et al. [2004] discuss an analytic solution for the average path length in a random graph, within which they provide a method to find the probability that the geodesic distance between two nodes is exactly $k$. Rastelli et al. [2016] uses the same reasoning and finds an explicit expression under the Gaussian latent position model(Gaussian LPM, GLPM), which is given by:
@@
To summarize this method: First, define $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as the probability of at least one walk of length $k$ between nodes $i$ and $j$, where $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ are their latent positions respectively. A walk is called a path only when all the nodes it passes through are distinct, while a walk does not require all its nodes to be distinct. In other words, $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ could also be thought of as the probability that the geodesic distance between these two nodes is not larger than $k$. Then, define $\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ as the probability that the geodesic distance between nodes $i$ and $j$ is exactly $k$. Given the definition of $p_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$, we have:
@@
We further define $\xi_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ to be the probability that there exits a walk of length $k$ from node $i$ to node $j$ that passes through nodes $\{n_{i},n_{A_{m}^{1}},\dots,n_{A_{m}^{k-1}},n_{j}\}$, and we refer to this walk of $k+1$ nodes as event $A_{m}$. By the definition of a walk, we do not require $n_{A_{m}^{1}},\dots,n_{A_{m}^{k-1}}$ to be distinct. The number of such events is about $N^{k-1}$. Let $A=\cup_{m=1}^{N^{k-1}}A_{m}$, the event that at least one walk of length $k$ exists between nodes $i$ and $j$. By Lemma 1 of Fronczak et al. [2004], we have:
@@
As suggested in Rastelli et al. [2016], for an observed network, assuming $\gamma=1$, $\phi$ and $\tau$ could be found via an ad-hoc method that matches the observed and theoretical values of average degree and the clustering coefficient through equation (15) and (16).
3.2 The conditional latent distance distribution
By definition, $\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})$ is a conditional probability, that is,
$\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|\mathbf{z}_{i},\mathbf{z}_{j}).$ (9)
Based on the Gaussian LPM (3), we know that the probability of having an edge depends only on $d_{ij}=d_{\mathbb{R}^{d}}(\mathbf{z}_{i},\mathbf{z}_{j})$. Thus, we could also rewrite equation (9) as
$\ell_{k}(\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|\mathbf{z}_{i},\mathbf{z}_{j})=P(\delta_{ij}=k|d_{ij}).$ (10)
Then, the distribution of the latent distance between nodes given their geodesic distance could be found through:
@@ 3.2 The conditional latent distance distribution
For the marginal distribution of latent distances, consider that when $d=2$ and $\gamma=1$,
@@
The Figure 2 shows the comparison between the conditional latent distribution obtained through (11) and the histogram of latent distances between nodes given geodesic distance of two networks. Both networks are generated using Gaussian LPM, and the size is 50. In general, the theoretical distribution is broadly consistent with the empirical histograms, especially when geodesic distances are small. As the geodesic distance increases, the theoretical distributions become increasingly concentrated and shift toward larger latent distances. In contrast, the empirical histograms show more spread and include a substantial mass at smaller latent distances, suggesting that the theoretical model over-concentrates probability on larger distances and underrepresents shorter latent distances for higher geodesic levels. However, because these longer geodesic distances make up only a small proportion of the full distribution, this discrepancy is unlikely to substantially affect overall performance.
(a) Network A, Density: 0.0988
(b) Network B, Density: 0.5176
Figure 2: Comparison between the theoretical conditional distribution of latent distances,shown by the curve, and the empirical histogram of observed latent distances between nodes. Both networks are of size 50 and generated using Gaussian LPM (3) with $\gamma=1$ and $\phi=2$, $\tau=0.2$ for Network A and $\tau=1$ for Network B.
3.3 Nonnested hypothesis tests
The $J$-test is a classical way to compare two nonnested regression models. The core idea is that if model $M_{2}$ captures additional structure missing from model $M_{1}$, then $M_{2}$’s fitted values should add explanatory power. In Section 5 of Davidson and MacKinnon [2002], they discuss the use of a parametric bootstrapped $J$ test, with hypotheses:
$\mathbf{H_{0}}$ :
The competing model $M_{2}$ does not provide additional explanatory power beyond the current model $M_{1}$.
$\mathbf{H_{A}}$ :
The competing model $M_{2}$ provides additional explanatory power.
To start with, they find the $J$ statistic for the observed data set, denoted by $\hat{J}$. Then, they form a bootstrap distribution of the $J$ statistic. They fit the data into model $M_{1}$, generate parameter estimates, and use them in the bootstrap data generating process. Once they have $B$ bootstrap samples, $J$ statistics are calculated again for these samples using the same method as the observed data set and are denoted by $J_{j}^{*},j=1,\dots,B$. Finally, the p-value for this test is given by
@@ 3.3 Nonnested hypothesis tests
In the case of the latent space model, we adopt this bootstrapping method of the Davidson-MacKinnon $J$ test. Similar to the Davidson-MacKinnon $J$-test, we are comparing two nonnested models: GLPM as current model and a hyperbolic latent position model as competing model. The hypotheses are:
$\mathbf{H_{0}}$ :
The hyperbolic model does not provide additional explanatory power beyond the euclidean model.
$\mathbf{H_{A}}$
The hyperbolic model provides additional explanatory power.
We first find the stress difference between classical and hyperbolic MDS. Then, just as they estimate model parameters to generate bootstrap data, we use the conditional distribution of the latent distance between nodes given geodesic distance under the GLPM to generate bootstrap networks under this model. Then, we calculate stress differences for these bootstrap networks and find the p-value. This method is described in Algorithm 2.
Algorithm 2 Bootstrapped Networks: Conditional latent-distance sampling
@@ 4 Simulations
We want to show the efficiency of the methods mentioned in the previous sections. To achieve this goal, we simulate networks in Euclidean and hyperbolic latent spaces, using the models described in the following sections.
@@ 4.1 Latent space models
Modeling assumptions for the latent space models are as follows:
@@ 4.1.1 Euclidean Geometry
Most commonly mentioned latent space models for networks that use Euclidean geometry as underlying latent space is the ”Distance Models” proposed in Hoff et al. [2002]. Later, Rastelli et al. [2016] proposed the Gaussian latent position model, in which a non-normalized Gaussian density is used for edge instead of logistic link function in Distance Models. We will use the Gaussian LPM (3) as the Euclidean latent space models. Rastelli et al. [2016] has also shown that both expected average degree and clustering coefficient have explicit forms for Gaussian LPM, and are given by: